# Quantum Convolutional Neural Networks for Jet Images Classification

### v2

@@ Abstract
Recently, interest in quantum computing has significantly increased, driven by its potential advantages over classical techniques. Quantum machine learning (QML) exemplifies one of the important quantum computing applications that are expected to surpass classical machine learning in a wide range of instances. This paper addresses the performance of QML in the context of high-energy physics (HEP). As an example, we focus on the top-quark tagging, for which classical convolutional neural networks (CNNs) have been effective but fall short in accuracy when dealing with highly energetic jet images. In this paper, we use a quantum convolutional neural network (QCNN) for this task and compare its performance with CNN using a classical noiseless simulator. We compare various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. For every quantum setup, we design a similar setup to the corresponding classical model for a fair comparison. Our results indicate that QCNN with proper setups tend to perform better than their CNN counterparts, particularly when the convolution block has a lower number of parameters. For the higher parameter regime, the QCNN circuit was adjusted according to the dimensional expressivity analysis (DEA) to lower the parameter count while preserving its optimal structure. The DEA circuit demonstrated improved results over the comparable classical CNN model.
@@ I Introduction
Classical convolutional neural networks (CNNs) have been widely and successfully employed for top-quark tagging [15]. However, they struggle to provide the required accuracy when faced with the highly energetic and complex top-quark jet images [16]. Aiming to overcome this problem, our main objective is to identify a practically suitable quantum machine learning architecture to classify top-quark and QCD jet images. We do this by quantifying its accuracy and comparing its performance with a classical machine learning method. Specifically, we focus on the quantum convolutional neural network (QCNN), first introduced by Cong et al. [17], which leverages shallow-depth quantum circuits, making it suitable for the noisy intermediate-scale quantum (NISQ) devices currently accessible. This circuit shallowness also provides robustness against barren plateau issues, ensuring the trainability of the model [18]. While these advantages might make QCNN classically simulable, as highlighted in [19], the classical simulation of QCNN can still be regarded as a quantum-inspired machine learning architecture. In fact, several previous works have demonstrated improved accuracy of QCNN over conventional classical CNNs under certain conditions [20, 21].
@@
The structure of this paper goes as follows. Section II describes the methodology of our study. In this section, we start by introducing the main components of a basic classical CNN model. We then explain how these concepts can be translated into the quantum domain to construct a QCNN, which is subsequently analyzed using DEA. After that, we discuss the formation and the pre-processing of the jet images. The results and QCNN to CNN comparisons are presented in section III. Then, we conclude and mention future directions and further plans of our study in section IV.
II Methodology
II.1 CNN
@@
II.2 Loss functions
@@
2) The Mean-Square-Error (MSE) considers binary labels ±$\pm 1$ and is given by
@@
3) The Hinge loss also takes ±$\pm 1$ labels as inputs and defined as
@@
II.3 QCNNs
QCNN was a QNN model motivated by the success of CNNs [17]. The basic idea of QCNN is to utilize quantum operations to accomplish the convolution and pooling techniques and do the classification task using the outcome from the circuit. In the case of classical inputs (like images), we need to encode the data (image pixels) first as |$\ket{\psi(\bm{x})}=U_{\text{enc}}(\bm{x})\ket{0}$, where $\bm{x}$ is classical data. Then the encoded state is processed via parametrized unitaries composed of convolutional and pooling blocks to have an output $\hat{y}$, given by
$\hat{y}=\braket{\psi(\bm{x})}{U^{{\dagger}}_{\text{QCNN}}(\bm{\theta})Z_{N_{q}}U_{\text{QCNN}}(\bm{\theta})}{\psi(\bm{x})}\,,$ (4)
@@ II.3 QCNNs
II.3.1 Encodings
The encoding $V_{\text{enc}}$ can be done in several ways. We test four encoding schemes, namely: tensor product embedding (TPE) which in this case is also known as angle encoding, hardware efficient embedding (HEE), and Classically Hard Embedding (CHE). These three encodings are used to study the effects of encoding on trainability in [27]. The quantum circuit structure of the different encodings can be seen in Fig. 3, where TPE is shown in Fig. 3a, HEE in 3b, and CHE in 3c. The data inputs in these figures are represented by $x_{i}$, where $i\in\{0,1,2,3\}$. In this paper, we use one layer and two layers of HEE, denoted by HEE1 and HEE2, respectively. On the other hand, only one layer of CHE is used in this work. Since reference [27] showed that increasing the number of layers would cause trainability issues, we only focus on these encodings with the limited number of layers.
@@ II.3.1 Encodings
II.3.2 Convolutional and Pooling circuits
After the encoding layer comes the convolutional layer. This layer consists of convolutional circuits connecting every pair of qubits to compensate for the task of a classical convolutional filter that scans over all the pixel values of an image. The convolutional circuit can be built in various ways. In this project, we test two unitary two-qubit circuits $SO(4)$ and $SU(4)$. The dimensionality of the trainable parameters $\bm{{\theta}}$ depends on these two-qubit unitaries. The circuit constructions for these unitaries are presented in Fig. 4a and 4b. The $SO(4)$ circuit corresponds to real transformation [28] and it contains 2 CNOTs and 12 single-qubit gates of which only 6 of them contain trainable parameters ($\bm{{\theta}}$). $SU(4)$ is more general and considered a universal two-qubit circuit spanning the whole Hilbert space. It consists of 3 CNOTs and 15 single-qubit gates all containing trainable parameters [28, 29]. Then, the convolutional circuit connects every pair of neighboring qubits until it forms a full convolutional layer.
After that comes the pooling layer to reduce the dimensionality of the system by half. The pooling circuit, shown in Fig. 4c, comprises a CNOT gate and 9 single-qubit gates all containing trainable parameters. The convolutional and pooling layers are repeated until the system has reached a sufficiently small size, one qubit in our case. The output of this qubit is measured with a PauliZ measurement. The prediction of the neural network corresponds to the expectation value of the PauliZ operator and is then compared to the true label for the classification task as explained in Section II.2.
@@ II.5 Dataset Preparation
II.5.1 Jet images
@@
In this project, $\gamma_{B}=10$ is chosen.
@@
This means that the components $X_{i}$ and $Y_{i}$ can only be in the range of $[-1,1]$. These components then form the coordinates of the two-dimensional histogram corresponding to the jet image. The histogram is filled with the weight
@@
II.5.2 PCA
Principal Component Analysis (PCA) (see e.g. [33]) is a method that is commonly used to find patterns while reducing the dimensionality of data. It mainly captures correlation in data and transforms it into a new set of variables that better describes the patterns. These variables are called the principal components. In this project, PCA is used to reduce the dimensionality of the jet images from $28\times 28$ to $2\times 2$, so that we end up with only 4 pixels that are the inputs of our QCNN and CNN models.
III Results
@@
[…]
III.1 Dependence on Encodings
@@
The accuracy curves corresponding to these settings are also summarized in Fig. 6 for $SO(4)$ and $SU(4)$ compared to CNN with Hinge (Fig. 6(a) and 6(b)), MSE (Fig. 6(c) and 6(d)), and Cross-entropy (Fig. 6(e) and 6(f)) loss functions. We can notice the faster convergence at early epochs for all setups of QCNN for both $SO(4)$ and $SU(4)$ compared to CNN.
@@
III.2 Dependence on Batch-sizes
[…]
$SO(4)$ H 16 99.56 ± 0.05 87.79 ± 3.09
32 99.48 ± 0.07 90.84 ± 2.78
64 99.05 ± 0.18 89.46 ± 2.87
128 96.48 ± 0.62 77.86 ± 3.41
M 16 99.90 ± 0.04 93.87 ± 2.35
32 99.86 ± 0.05 89.76 ± 2.89
64 99.70 ± 0.08 93.72 ± 2.35
128 98.23 ± 0.25 91.50 ± 2.38
C 16 99.71 ± 0.05 91.88 ± 2.66
32 99.55 ± 0.06 91.78 ± 2.65
64 99.28 ± 0.17 93.59 ± 2.34
128 97.53 ± 0.30 98.23 ± 1.03
$SU(4)$ H 16 99.43 ± 0.08 99.97 ± 0.02
32 99.47 ± 0.07 99.95 ± 0.03
64 99.14 ± 0.12 97.87 ± 1.42
128 98.52 ± 0.22 94.81 ± 1.82
M 16 99.98 ± 0.01 100.00 ± 0.00
32 99.95 ± 0.04 99.96 ± 0.02
64 99.84 ± 0.04 99.92 ± 0.03
128 99.43 ± 0.13 99.36 ± 0.22
C 16 99.54 ± 0.06 100.00 ± 0.00
32 99.60 ± 0.07 99.995 ± 0.005
64 99.40 ± 0.10 98.76 ± 1.01
128 98.93 ± 0.17 98.77 ± 0.41
@@ III.2 Dependence on Batch-sizes
After fixing the encoding to HEE1 which shows the best performance for $32$ batch-size, we varied the batch size to four different sizes. The results averaged over $50$ runs are displayed in Table 2. Again we observe a tendency for QCNN to have a better performance for most of the batch-sizes in $SO(4)$ compared to CNN, and results are comparable for $SU(4)$ and CNN. However, what is clearly noticeable in most of the cases, is that as the batch size increases, the accuracy tends to decrease. This observation is also depicted in Fig. 7 for $SU(4)$ and the corresponding CNN model with $51$ parameters for Hinge loss as an example. We also observe that a batch size of $16$ has faster convergence compared to the other batch sizes, while a batch size of $128$ exhibits the slowest convergence. Batch sizes of 16 and 32 demonstrate similar performance (values shown in bold in Table 2), which is superior to both 64 and 128. It is essential to note that slower or faster convergence does not necessarily equate to slower or faster training times, since the time it takes to train each epoch is different from quantum to classical.
III.3 DEA circuit

### v3

@@ Abstract
Recently, interest in quantum computing has significantly increased, driven by its potential advantages over classical techniques. Quantum machine learning (QML) exemplifies one of the important quantum computing applications that are expected to surpass classical machine learning in a wide range of instances. This paper addresses the performance of QML in the context of high-energy physics (HEP). As an example, we focus on the top-quark tagging, for which classical convolutional neural networks (CNNs) have been effective but fall short in accuracy when dealing with highly energetic jet images. In this paper, we use a quantum convolutional neural network (QCNN) for this task and compare its performance with CNN using a classical noiseless simulator. We compare various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. For every quantum setup, we design a similar setup to the corresponding classical model for a fair comparison. Our results indicate that QCNN with proper setups tend to perform better than their CNN counterparts, especially when the convolution block has a lower number of parameters. For the higher parameter regime, the QCNN circuit was adjusted according to the dimensional expressivity analysis (DEA) to lower the parameter count while preserving its optimal structure. The DEA circuit demonstrated improved results over the comparable classical CNN model.
@@ I Introduction
Classical convolutional neural networks (CNNs) have been widely and successfully employed for top-quark tagging [15]. However, they struggle to provide the required accuracy when faced with the highly energetic and complex top-quark jet images. The complexity of an image is defined by how dense the information it holds and how diverse its visual features are [16]. Therefore, many parameters will be needed to extract such features from the images which is further discussed in Section III.1. In [17], various modern machine learning models, including CNN as an image-based model, were benchmarked for the top-tagging task. These models demonstrated comparative performance, with accuracies ranging from 90% to 94%, but required intensive trainable parameter counts ranging from 1000 to 1.46 millions. The latter parameter count corresponds to the state-of-the-art CNN-based architecture named ResNeXt. This architecture introduces a balance between model depth and width for efficient feature learning, however, at the cost of increased computational cost and memory usage [18]. Aiming to overcome this challenge of increased parameter count, our main objective is to identify a practically suitable QML architecture to efficiently classify top-quark and QCD jet images. We do this by quantifying its accuracy and comparing its performance with a classical machine learning method. Specifically, we focus on the quantum convolutional neural network (QCNN), first introduced by Cong et al. [19], which leverages shallow-depth quantum circuits, making it suitable for the noisy intermediate-scale quantum (NISQ) devices currently accessible. This circuit shallowness also provides robustness against barren plateau issues, ensuring the trainability of the model [20]. While these advantages might make QCNN classically simulable, as highlighted in [21], the classical simulation of QCNN can still be regarded as a quantum-inspired machine learning architecture. In fact, several previous works have demonstrated improved accuracy of QCNN over conventional classical CNNs under certain conditions [22, 23].
@@
The structure of this paper is as follows. In Section II, the concept of boosted top-quarks originating from a $W^{\prime}$ boson is presented. Section III describes the methodology of our study. In this section, we start by introducing the main components of a basic classical CNN model. We then explain how these concepts can be translated into the quantum domain to construct a QCNN, which is subsequently analyzed using DEA. After that, we discuss the formation and the pre-processing of the jet images. The results and QCNN to CNN comparisons are presented in section IV. Then, we conclude and mention future directions and plans in section V.
II High-Energy Physics Background
The Standard Model of particle physics successfully describes the fundamental particles and three of the four fundamental forces, excluding gravity, governing their interactions. Despite its success, it is incomplete, as it does not explain phenomena like gravity, dark matter, or neutrino mass. Extensions beyond the Standard Model (BSM) aim to address these gaps, and searches for BSM particles are among central objectives at the LHC experiments.
A leading candidate among the hypothetical particles proposed by various BSM theories is the $W^{\prime}$ boson. This particle is considered a heavier analog of the Standard Model $W$ boson and is theorized to interact predominantly with third-generation fermions [27], such as the top quark. A specific type, the right-handed $W^{\prime}$ boson, is predicted within the left-right symmetric model, which is an extension of the Standard Model’s electroweak sector that unifies the gauge groups as $SU(2)_{R}\times SU(2)_{L}\times U(1)$ [28].
The D0 Collaboration initiated one of the earliest searches for the $W^{\prime}$ boson in 1995, focusing on its decay into an electron and an anti-neutrino. This study established exclusion limits on the $W^{\prime}$ boson mass to 800GeV [28, 29]. Subsequent analyses, such as a 2017 study by the CMS Collaboration, placed a 95% confidence level lower bound of 2.4TeV on the mass of a right-handed $W^{\prime}$ boson decaying into a top and a bottom quark [30].
Because of its large predicted mass, the $W^{\prime}$ boson is expected to decay promptly into other particles. One of the most studied decay channels is $W^{\prime}\rightarrow tb$, which is especially compelling due to its relatively low background from QCD processes [30]. Figure 1 depicts the leptonic decay of the $W^{\prime}$ via this mode. Thus, developing an efficient and reliable approach for tagging boosted top-quark jets could provide deeper insights into the properties of the $W^{\prime}$ boson in case it was produced in a particle collider.
Figure 1: Leptonic decay channel of $W^{\prime}$ boson to a top and a bottom quark ($W^{\prime}\rightarrow tb$).
Top quark identification in collider experiments involves detecting characteristic features such as secondary vertices from $b$-quark decays [27] and localized energy deposits in detectors corresponding to jets formed by the final-state particles. The neutrino, which escapes direct detection, is inferred from the missing transverse energy.
Several methods are employed to analyze these decays. One approach involves creating a two-dimensional histogram of the energy distribution of decay products, typically a lepton and $b$-quark hadrons, which are represented as two subjets in the resulting jet image. The formation of these jet images is described in detail in Section III.5.2. The angular separation between the subjets is determined by the transverse momentum ($p_{T}$) of the top quark. When the top quark is highly boosted, this angle becomes small, leading to an overlap between the lepton and $b$-quark subjets. This phenomenon, illustrated in Figure 2, shows how boosted top jets, with high $p_{T}$, have closely aligned subjets, while non-boosted (low $p_{T}$) tops exhibit well-separated ones. This overlap can cause boosted top jet images to resemble those from QCD processes.
Figure 2: Illustration of boosted jets, e.g. decayed from a top quark.
III Methodology
III.1 CNN
@@
A detailed structure of the CNN model used in this study, consisting of 33 trainable parameters, is presented in Fig. 4. The first layer, labeled Conv2d, contains four feature map kernels, each designed to extract specific patterns from the input image. Generally, as the complexity of input images increases, a larger number of feature maps are required to capture more features, which in turn leads to a higher number of trainable parameters. In other words, complex datasets require complex models that are capable of learning complex connections and patterns found in data points. This can significantly increase the computational cost, potentially making the model impractical for large or complex datasets. Such models also suffer from the overfitting issue which results from the increased number of trainable parameters which perfectly fits to the given dataset, but cannot generalize to new dataset [32]. This challenge represents a key limitation of conventional CNNs and traditional machine learning models. For example, as previously mentioned in the introduction, ResNeXt representing a state-of-the-art CNN architecture involves number of parameters exceeding a million. Methods such as regularization and principal component analysis (PCA) are usually employed to decrease the complexity and increase generalization of machine learning models. The used case of this project utilizes PCA to reduce the dimensionality of the dataset which reduces the number of parameters. However, this might cause a reduction in the efficiency of the classical machine learning model.
One of the goals of QML models is to address this issue by employing quantum entanglement, which allows the model to learn correlations between features across the image rather than relying on local information only. The main subject of this paper, QCNNs, are an example of such models and are discussed in Section III.3.
Figure 4: Detailed structure of the CNN model used in this study.
III.2 Loss functions
@@
2) The Mean-Square-Error (MSE) considers binary labels $\pm 1$ and is given by
@@
3) The Hinge loss also takes $\pm 1$ labels as inputs and defined as
@@
III.3 QCNNs
[…]
$\hat{y}=\braket{\psi(\bm{x})|U^{{\dagger}}_{\text{QCNN}}(\bm{\theta})Z_{N_{q}}U_{\text{QCNN}}(\bm{\theta})|\psi(\bm{x})}\,$, (4)
@@ III.3 QCNNs
III.3.1 Encodings
The encoding $V_{\text{enc}}$ can be done in several ways. We test four encoding schemes, namely: tensor product embedding (TPE) which in this case is also known as angle encoding, hardware efficient embedding (HEE), and Classically Hard Embedding (CHE). These three encodings are used to study the effects of encoding on trainability in [34]. The quantum circuit structure of the different encodings can be seen in Fig. 6, where TPE is shown in Fig. 6a, HEE in 6b, and CHE in 6c. The data inputs in these figures are represented by $x_{i}$, where $i\in\{0,1,2,3\}$. In this paper, we use one layer and two layers of HEE, denoted by HEE1 and HEE2, respectively. On the other hand, only one layer of CHE is used in this work. Since reference [34] showed that increasing the number of layers would cause trainability issues, we only focus on these encodings with the limited number of layers.
@@ III.3.1 Encodings
III.3.2 Convolutional and Pooling circuits
After the encoding layer comes the convolutional layer. This layer consists of convolutional circuits connecting every pair of qubits to compensate for the task of a classical convolutional filter that scans over all the pixel values of an image. The convolutional circuit can be built in various ways. In this project, we test two unitary two-qubit circuits $SO(4)$ and $SU(4)$. The dimensionality of the trainable parameters $\bm{{\theta}}$ depends on these two-qubit unitaries. The circuit constructions for these unitaries are presented in Fig. 7a and 7b. The $SO(4)$ circuit corresponds to real transformation [35] and it contains 2 CNOTs and 12 single-qubit gates of which only 6 of them contain trainable parameters ($\bm{{\theta}}$). $SU(4)$ is more general and considered a universal two-qubit circuit spanning the whole Hilbert space. It consists of 3 CNOTs and 15 single-qubit gates all containing trainable parameters [35, 36]. Then, the convolutional circuit connects every pair of neighboring qubits until it forms a full convolutional layer.
After that comes the pooling layer to reduce the dimensionality of the system by half. The pooling circuit, shown in Fig. 7c, comprises a CNOT gate and 9 single-qubit gates all containing trainable parameters. The convolutional and pooling layers are repeated until the system has reached a sufficiently small size, one qubit in our case. The output of this qubit is measured with a PauliZ measurement. The prediction of the neural network corresponds to the expectation value of the PauliZ operator and is then compared to the true label for the classification task as explained in Section III.2.
@@ III.5 Dataset Preparation
III.5.1 JetNet Dataset
JetNet is a Python library developed to enhance accessibility and reproducibility in machine learning research within HEP, with a main focus on particle jet analysis. It is built on PyTorch framework and provides interfaces for accessing a range of HEP datasets, along with built-in evaluation metrics and general-purpose tools tailored for HEP tasks [24].
The library includes three main dataset categories: JetNet, QuarkGluon, and TopTagging. The JetNet dataset contains jets initiated by gluons, top quarks, light quarks, as well as W and Z bosons. As its name suggests, the QuarkGluon dataset consists of jets originating from quarks and gluons. The TopTagging dataset, which is used in the analysis presented in this study, features hadronic top jets as signal and QCD jets as background. All datasets are based on Monte Carlo simulations, which model particle collisions and detector responses to replicate real experimental conditions [39].
In addition, the JetNet dataset offers functionality for visualizing jet images using histogram plots. Some of these examples are shown in Fig. 8. Each jet image corresponds to an individual jet from a separate event, mapped onto the $\eta$$\phi$ plane. The intensity of each pixel reflects the transverse momentum of the jet’s constituents at that location. These individual images usually do not reveal any clear structure or consistent pattern within a given class. To uncover broader features, particle physicists often sum and normalize thousands of such jet images, creating an average or composite image that emphasizes recurring patterns not visible in single-event displays, which will be presented in the next section.
Figure 8: Jet images of a single-jet each of gluon (g) and top (t) particles [24].
III.5.2 Jet images
@@
In this project, $\gamma_{B}=10$ is chosen which produced a convenient jet shape compared to several values arbitrarily tested.
@@
This means that the components $X_{i}$ and $Y_{i}$ can only be in the range of [1,1][-1,1]. These components then form the coordinates of the two-dimensional histogram corresponding to the jet image. The histogram is filled with the weight
@@
III.5.3 PCA
Principal Component Analysis (PCA) (see e.g. [41]) is a method that is commonly used to find patterns while reducing the dimensionality of data. It mainly captures correlation in data and transforms it into a new set of variables that better describes the patterns. These variables are called the principal components.
In this study, PCA is used to reduce the dimensionality of the jet images from $28\times 28$ to $2\times 2$, so that we end up with only 4 pixels that are the inputs of our QCNN and CNN models. This implies that the original pixel data are projected onto the four principal eigenvectors of the data’s covariance matrix, which are those associated with the largest eigenvalues. These principal components preserve most of the essential information from the original images, providing a compact and efficient representation of the input. Representative examples of such projected images for both top and QCD jet classes are shown in Fig. 10.
Figure 10: Jet images of top-quark (Class 1) and QCD (Class 0) originally with size 28x28 (top) compared to the PCA reduced of 2x2 (bottom).
IV Results
@@
$\text{(accuracy)}:=\frac{\text{(correct classifications)}}{\text{(the number of total samples)}}\,.$ (16)
IV.1 Dependence on Encodings
@@
The accuracy curves corresponding to these settings are also summarized in Fig. 11 for $SO(4)$ and $SU(4)$ compared to CNN with Hinge (Fig. 11(a) and 11(b)), MSE (Fig. 11(c) and 11(d)), and Cross-entropy (Fig. 11(e) and 11(f)) loss functions. We can notice the faster convergence at early epochs for all setups of QCNN for both $SO(4)$ and $SU(4)$ compared to CNN.
@@
IV.2 Dependence on Batch-sizes
Circuit Loss Batch-size QCNN Acc ($\%$) CNN Acc ($\%$)
$SO(4)$ H 16 99.93 ± 0.07 94.70 ± 1.74
32 99.73 ± 0.12 92.48 ± 1.79
64 99.23 ± 0.23 89.94 ± 2.02
128 97.48 ± 0.82 69.30 ± 2.90
M 16 100.00 ± 0.00 94.12 ± 2.00
32 99.94 ± 0.03 94.76 ± 1.84
64 99.69 ± 0.10 87.38 ± 2.48
128 98.89 ± 0.34 87.35 ± 1.33
C 16 100.00 ± 0.00 92.48 ± 2.30
32 99.91 ± 0.05 92.93 ± 1.92
64 99.43 ± 0.16 88.54 ± 2.47
128 98.23 ± 0.54 83.38 ± 1.97
$SU(4)$ H 16 99.75 ± 0.14 97.95 ± 0.44
32 99.92 ± 0.08 97.56 ± 0.45
64 99.72 ± 0.11 93.36 ± 1.63
128 99.22 ± 0.17 85.35 ± 2.06
M 16 99.84 ± 0.09 98.80 ± 0.29
32 99.94 ± 0.05 98.32 ± 0.37
64 99.86 ± 0.06 97.02 ± 0.46
128 99.64 ± 0.09 90.88 ± 1.34
C 16 99.82 ± 0.10 97.16 ± 1.13
32 99.90 ± 0.07 96.63 ± 1.26
64 99.79 ± 0.08 95.86 ± 0.68
128 99.35 ± 0.14 91.57 ± 0.89
@@ IV.2 Dependence on Batch-sizes
After fixing the encoding to HEE1 which shows the best performance for $32$ batch-size, we varied the batch size to four different sizes. The results averaged over $50$ runs are displayed in Table 2. Again we observe a tendency for QCNN to have a better performance for all the batch-sizes in $SO(4)$ compared to CNN, and results are comparable for $SU(4)$ and CNN. However, what is clearly noticeable in most of the cases, is that as the batch size increases, the accuracy tends to decrease. This observation is also depicted in Fig. 12 for $SU(4)$ and the corresponding CNN model with $51$ parameters for Hinge loss as an example. We also observe that a batch size of $16$ has faster convergence compared to the other batch sizes, while a batch size of $128$ exhibits the slowest convergence. Batch sizes of 16 and 32 demonstrate similar performance (values shown in bold in Table 2), which is superior to both 64 and 128. It is essential to note that slower or faster convergence does not necessarily equate to slower or faster training times, since the time it takes to train each epoch is different from quantum to classical.
IV.3 DEA circuit